The remainder of the division of the polynomial x2 +1 by the polynomial x-1 is
A. 0
B. 1
C. 2
D. None of the above
1
Expert's answer
2021-08-23T17:18:04-0400

"\frac{x^2+1}{x-1}=[x+1+\frac{2}{x-1}]"


Hence the remainder = 2